# anaghmalik.com/FlyingWithPhotons

### v2

@@ 2 Related Work
Our work is closely related to time-resolved imaging modalities that directly capture light in flight [62, 44, 30]. Among these modalities, interferometry provides the highest temporal resolution and resolves light propagation at micron scales [13, 25]. Still, interferometric techniques typically require bulky optical setups and have a limited working range, and so are more suited to microscopic imaging. Streak cameras achieve sub-picosecond resolution [66], but are also bulky and expensive, making them more difficult to use. On the other hand, photonic mixer devices are far less expensive and can be used to reconstruct transient videos, though at coarser, nanosecond scales [15, 45]. SPADs provide a middle ground between these sensing modalities, as they are relatively inexpensive, yet have high temporal resolution at the picosecond level [52]. Our work uses a scanned, single-pixel SPAD with $\approx$50 ps temporal resolution.
Various other methods have been developed to capture phenomena relating to light transport. For example, structured illumination can separate light into different components (e.g., direct or indirect) by projecting high frequency patterns [41, 14], illuminating and imaging based on epipolar lines [46, 43], or using a combination of structured illumination and masking [47]. Other work captures light paths directly by imaging scenes placed into a fluorescent medium [16]. However, none of these techniques can capture propagating light.
@@
Our work builds on methods for 3D reconstruction and novel view synthesis based on neural radiance fields (NeRFs) [37]. While such techniques have been developed to, e.g., model refractive objects [3, 48, 72] and glossy surfaces [63], or to perform inverse rendering tasks [74, 33, 22, 24, 38, 56, 58, 64], no previous method performs novel view synthesis of complex, time-resolved light transport effects observed from multiple viewpoints. To do so requires a dataset that captures these effects and a representation to facilitate rendering them. Neither of which are currently available; our method provides both.
Closest to our work is TransientNeRF [34], which performs novel view synthesis of the direct component of transient measurements i.e. from the data used in lidar. However, TransientNeRF cannot render indirect light transport effects, and relies on a lidar-specific image formation model, where the sensor and light source are coaxial. In contrast our approach can handle a non-coaxial light source, such as a point source or collimated beam placed anywhere within a scene. Further, we can render more general light transport effects, such as diffuse and specular reflections, multiple scattering, refraction, and diffraction.
@@ 3.2 Transient Fields and Rendering
We use a volumetric representation of the scene to render the transient $\boldsymbol{\tau}_{\mathbf{r}}\in\mathbb{R}_{+}^{N}$ given a camera ray $\mathbf{r}$ (see Fig. 2). Specifically, we define a point on a ray $\mathbf{r}(s)=\mathbf{o}+s\mathbf{d}$ where $\mathbf{o}\in\mathbb{R}^{3}$ is the camera center of projection and $\mathbf{d}$ is a three-dimensional unit vector corresponding to the ray direction. The transient field, $\boldsymbol{\tau}_{\theta}:(\mathbf{r}(s),\mathbf{d})\mapsto\mathbb{R}_{+}^{N}$, is a neural representation with parameters $\theta$ that maps a three-dimensional point and a ray direction to the discrete time bin values of a transient. Additionally, the neural representation outputs a spatially varying absorption coefficient $\sigma(\mathbf{r}(s))$, which gives the differential probability of ray termination at each point in the volume. Finally, the transient along ray $\mathbf{r}$ is computed using a modified version of the volume rendering equation:
@@
Here, $T$ represents transmittance from ray distance $s_{n}$ to $s$, and $c$ is the speed of light. The transient vectors $\boldsymbol{\tau}_{\theta}$ sampled along the ray are time shifted via convolution ($\ast$) with the Kronecker delta function $\delta$—this accounts for the propagation delay to the camera and is similar to the time warping procedure of Velten et al. [62]. If the time shift is not modeled, there is an ambiguous mapping from $(\mathbf{r}(s),\mathbf{d})$ to shifted versions of the same transient, depending on the distance to the camera center of projection. The volume rendering integral is evaluated numerically using the quadrature rule proposed by Max [36].
@@ 4 Multi-Viewpoint Transient Dataset
We use the renderer of Liu et al. [31] as well as a version of Mitsuba 2 modified for transient rendering [54] to simulate transient videos from multiple viewpoints. The illumination is a point source that emits a temporal impulse into the scene. For the Mitsuba 2 scenes we move the camera to sample a grid of 31 by 3 viewpoints on a hemisphere spanning 45 to 180 degrees in azimuth and 30 to 45 degrees in elevation. We create a total of 4 synthetic scenes based on the Cornell box or assets from Blendswap (https://blendswap.com/) and Bitterli [4], which we assemble together in Blender [59]. To evaluate the method, we render a set of 30 unseen viewpoints sampled on the same hemisphere as the training views, which is consistent with what we capture using our hardware system. We also capture another set of 30 unseen viewpoints that extend beyond the training views by sampling on a hemisphere of radius roughly $20\%$ greater than the training one. The simulated transients are rendered at 512x512 resolution and the number of time bins ranges from 300 to 900 depending on the scene scale.
@@
The captured dataset consists of 5 scenes, and for each scene we capture 45 (or 75) transient videos on a grid of 15 (or 25) by 3 viewpoints spanning 125 to 360 degrees in azimuth and 15 degrees in elevation. Capturing each transient video requires 20 to 30 minutes; we bin all captured photons into a transient histogram with 4096 bins, each spanning 4 ps. See the supplement for a detailed description of capture parameters and measured photon counts for each scene. Prior to using the simulated and captured datasets for view synthesis with our method, we normalize them as described in the supplement to ensure that the dynamic range is compatible with the output range of the neural representation. We will make all code and datasets publicly available upon publication.

### v3

@@ 2 Related Work
Our work is closely related to time-resolved imaging modalities that directly capture light in flight [64, 48, 33]. Among these modalities, interferometry provides the highest temporal resolution and resolves light propagation at micron scales [15, 28]. Still, interferometric techniques typically require bulky optical setups and have a limited working range. Streak cameras achieve sub-picosecond resolution [67], but are also bulky and expensive, making them more difficult to use. On the other hand, photonic mixer devices are far less expensive and can be used to reconstruct transient videos, though at coarser, nanosecond scales [18, 49]. SPADs provide a middle ground between these sensing modalities, as they are relatively inexpensive, yet have high temporal resolution at the picosecond level [56]. Our work uses a scanned, single-pixel SPAD with $\approx$50 ps temporal resolution.
Various other methods have been developed to capture phenomena relating to light transport. For example, structured illumination can separate light into different components (e.g., direct or indirect) by projecting high frequency patterns [45, 17, 41], illuminating and imaging based on epipolar lines [50, 47], or using a combination of structured illumination and masking [51]. Other work captures light paths directly by imaging scenes placed into a fluorescent medium [19]. However, none of these techniques can capture propagating light.
@@
Our work builds on methods for 3D reconstruction and novel view synthesis based on neural radiance fields (NeRFs) [40]. While such techniques have been developed to, e.g., model refractive objects [4, 52, 73] and glossy surfaces [65], or to perform inverse rendering tasks [75, 36, 25, 27, 42, 60], no previous method performs novel view synthesis of complex, time-resolved light transport effects observed from multiple viewpoints. To do so requires a dataset that captures these effects and a representation to facilitate rendering them. Neither of which are currently available; our method provides both.
Closest to our work is TransientNeRF [37], which performs novel view synthesis of the direct component of transient measurements, i.e., from the data used in lidar. However, TransientNeRF cannot render indirect light transport effects, and relies on a lidar-specific image formation model, where the sensor and light source are coaxial. In contrast, our approach handles non-coaxial light sources, such as a point source or collimated beam placed anywhere within a scene. Further, we render more general light transport effects, such as diffuse and specular reflections, multiple scattering, refraction, and diffraction.
@@ 3.2 Transient Fields and Rendering
We use a volumetric representation of the scene to render the transient $\boldsymbol{\tau}_{\mathbf{r}}\in\mathbb{R}_{+}^{N}$ given a camera ray $\mathbf{r}$ (see Fig. 2). Specifically, we define a point on a ray $\mathbf{r}(s)=\mathbf{o}+s\mathbf{d}$ where $\mathbf{o}\in\mathbb{R}^{3}$ is the camera center of projection and $\mathbf{d}$ is a three-dimensional unit vector corresponding to the ray direction. The transient field, $\boldsymbol{\tau}_{\theta}:(\mathbf{r}(s),\mathbf{d})\mapsto\mathbb{R}_{+}^{N}$, is a neural representation with parameters $\theta$ that maps a three-dimensional point and a ray direction to the discrete time bin values of a transient. Additionally, the neural representation outputs a spatially varying absorption coefficient $\sigma(\mathbf{r}(s))$, which gives the differential probability of ray termination at each point in the volume. Finally, the transient along ray $\mathbf{r}$ is computed using a modified version of the volume rendering equation [3, 16]:
@@
Here, $T$ represents transmittance from ray distance $s_{n}$ to $s$, and $c$ is the speed of light. The transient vectors $\boldsymbol{\tau}_{\theta}$ sampled along the ray are time shifted via convolution ($\ast$) with the Kronecker delta function $\delta$—this accounts for the propagation delay to the camera and is similar to the time warping procedure of Velten et al. [64]. If the time shift is not modeled, there is an ambiguous mapping from $(\mathbf{r}(s),\mathbf{d})$ to shifted versions of the same transient, depending on the distance to the camera center of projection. In practice we apply a continuous shift using linear interpolation. The volume rendering integral is evaluated numerically using the quadrature rule proposed by Max [39].
@@ 4 Multi-Viewpoint Transient Dataset
We use the renderer of Liu et al. [34] as well as a version of Mitsuba 2 modified for transient rendering [58] to simulate transient videos from multiple viewpoints. The illumination is a point source that emits a temporal impulse into the scene. For the Mitsuba 2 scenes we move the camera to sample a grid of 31 by 3 viewpoints on a hemisphere spanning 45 to 180 degrees in azimuth and 30 to 45 degrees in elevation. We create a total of 4 synthetic scenes based on the Cornell box or assets from Blendswap (https://blendswap.com/) and Bitterli [5], which we assemble together in Blender [6]. To evaluate the method, we render a set of 30 unseen viewpoints sampled on the same hemisphere as the training views, which is consistent with what we capture using our hardware system. We also render 30 additional unseen viewpoints by sampling on a hemisphere whose radius is roughly $20\%$ greater than the one used for training. The simulated transients are rendered at 512$\times$512 resolution and the number of time bins ranges from 300 to 900 depending on the scene.
@@
The captured dataset consists of 5 scenes, and for each scene we capture 45 (or 75) transient videos on a grid of 15 (or 25) by 3 viewpoints spanning 125 to 360 degrees in azimuth and 15 degrees in elevation. Capturing each transient video requires 20 to 30 minutes; we bin all captured photons into a transient histogram with 4096 bins, each spanning 4 ps. See the supplement for a detailed description of capture parameters and measured photon counts for each scene. Prior to using the simulated and captured datasets for view synthesis with our method, we normalize them as described in the supplement to ensure that the dynamic range is compatible with the output range of the neural representation. All code and datasets are publicly available on the project webpage.
@@ 6 Discussion
Acknowledgments
DBL and KNK acknowledge support of NSERC under the RGPIN, RTI, and Alliance programs. DBL also acknowledges support from the Canada Foundation for Innovation and the Ontario Research Fund. RP acknowledges the support of the Stanford Graduate Fellowship program.